OpenStudy (mrbee):

What are the potential solutions to the equation below? 2In(x+3)=0 They can be: x= -3 and x=-4 x= -2 and x=-4 x= 2 and x= -3 x= 2 and x= 4

OpenStudy (zzr0ck3r):

\(a\log(x+b)=0\implies \log(x+b)=0\implies 10^0=x+b\implies x=1-b\)

OpenStudy (welshfella):

ln( x + 3) = 0 what number has a logarithm of 0?

OpenStudy (mrbee):

Thank I got the answer x= -2 and x=-4

OpenStudy (welshfella):

yes

OpenStudy (zzr0ck3r):

why do they ask about potential solutions?

OpenStudy (zzr0ck3r):

So they want you to bring the exponent up, solve the quadratic and then exclude a solution. You can get around all of that by dividing the 2 away. Silly question.